OpenStudy (anonymous):

ANY1 HELP ME PLEASE The ratio of the radii of two circles at the centers of which two arcs of the same length subtended a angles of 60 degree and 70 degree is

OpenStudy (watchmath):

arc 1 =\(60/360\cdot 2\pi r_1\) arc 2 = \(70/360\cdot 2\pi r_2\) Since arc 1=arc 2 if we divide the first by the second we have \(1=6/7\cdot r_1/r_2\) So \(r_1/r_2=7/6\) where r1 is the outer radius and r2 is the inner radius

OpenStudy (anonymous):

YES BUT IT IS NOT THE OPTION SO I SHOULD THINK THAT THE QUESTION IS WRONG

OpenStudy (watchmath):

what are the options?

OpenStudy (anonymous):

A) 5:4 B) 4:5 C) 2:6 D) 4:3

myininaya (myininaya):

ok i don't mean to sound like a retard but can you tell me what subtended means

OpenStudy (anonymous):

extend opposite angle: to be opposite and delimit the extent of an angle or side of a geometric figure was not found in the English Dictionary .

OpenStudy (anonymous):

Subtended is a code word to let you know they are referring to a sector of a circle; they would usually say the subtended angle; referring to the sector bounded by radii, arc and subtended angle.